OpenStudy (anonymous):

Evaluate (1/100000)^3/4

OpenStudy (anonymous):

100000=10^5

OpenStudy (anonymous):

(1/10^5)^3/4 (1/10)^(5*3/4) (1/10)^(15/4)

OpenStudy (anonymous):

But I guess there its (1/10000)^3/4 not (100000)^3/4

OpenStudy (radar):

1/100000 same as\[10^{-5}\]

OpenStudy (radar):

\[(10^{-5})^{3/4}\]

OpenStudy (radar):

\[10^{(-5*3/4)}=10^{-15/4}= 10^{-3.75}\]

OpenStudy (radar):

Was this a multiple choice?